# Particle Semi-Implicit Variational Inference

### v1

@@ 3.1 Free energy functional
$\min_{(\theta,r)\in\mathcal{M}}\mathcal{E}(\theta,r),\quad\text{where}\enskip\mathcal{E}:\mathcal{M}\rightarrow\mathbb{R}:(\theta,r)\mapsto\int q_{\theta,r}(x)\log\frac{q_{\theta,r}(x)}{p(x,y)}\,\mathrm{d}x.$ (3)
@@
Assume that the evidence is bounded $|\log p(y)|<\infty$ and $k$ is bounded; then we have that $\mathcal{E}$ is (i) lower bounded, (ii) lower semi-continuous (l.s.c.), and (iii) non-coercive.
The proof can be found in Section E.1. Prop. 2 tells us that even though $\mathcal{E}$ possesses many of the properties one looks for in a meaningful minimization functional, it lacks coercivity (in the sense of Dal Maso (2012, Definition 1.12)), a sufficient property to establish the existence of solutions. The key to showing non-coercivity is that we can construct a kernel $k{{}_{\theta}}(x|z)$ that does not depend on $z$. At first glance, this issue might seem contrived but we note that this problem is closely related to the problem of posterior collapse (Lucas et al., 2019; Wang et al., 2021). To address non-coercivity, we propose to utilize regularization and define the regularized free energy as:
@@
Under the assumptions of Prop. 2, if $\mathsf{R}_{\lambda}$ is coercive and l.s.c. then, the regularized free energy $\mathcal{E}_{\lambda}$ is (i) lower bounded, (ii) l.s.c., (iii) coercive. Hence it admits at least one minimizer in $\mathcal{M}$.
The proof can be found in Section E.2. From here forward, we shall focus on the regularizer of the form $\mathsf{R}^{\mathrm{E}}_{\lambda}:(\theta,r)\mapsto\lambda_{r}\mathsf{KL}(r|p_{0})+\lambda_{\theta}\mathsf{R}_{\theta}(\theta)$ where $\lambda=\{\lambda_{r},\lambda_{\theta}\}$ are the regularization parameters and $p_{0}$ is a predefined reference distribution. There are many possible choices for $p_{0}$ and $\mathsf{R}_{\theta}$. For $p_{0}$, this regularizes solutions of the gradient flow toward it; as such, in settings where there is some knowledge or preference about $r$ at hand, we can set $p_{0}$ to reflect that. In our experiments, we utilize $p_{0}=\mathcal{N}(0,M)$ where $M$ is a positive definite (p.d.) matrix. As for $\mathsf{R}_{\theta}$, there are also many choices. In the context of neural networks, one natural choice is Tikhonov regularization $\frac{1}{2}\|\cdot\|^{2}$, resulting in weight decay (Hanson and Pratt, 1988) for gradient descent (Loshchilov and Hutter, 2017) which is a popular method for regularizing neural networks. In our experiments, we either use Tikhonov regularization or its simple variant $\|\theta\|^{2}_{M}:=\left\langle\theta,M\theta\right\rangle$. As long as $\mathsf{R}_{\theta}$ is coercive and $\lambda>0$, the resulting regularizor $\mathsf{R}^{\mathrm{E}}_{\lambda}$ will also be coercive.
@@ 3.2 Gradient flow
[…]
@@
$(\dot{\theta}_{t},\dot{r}_{t})=\nabla_{\mathcal{M}}\mathcal{E}_{\lambda}(\theta,r)\iff\begin{array}[]{l}\dot{\theta}_{t}=-\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\\ \dot{r}_{t}=-\nabla_{r}\mathcal{E}_{\lambda}(\theta_{t},r_{t})=\nabla_{z}\cdot\left(r_{t}\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]\right)\end{array}$ (7)
We now establish that the above gradient flow dynamics are contractive and that if a log-Sobolev inequality Eq. 9 holds, one can also establish exponential convergence. The log-Sobolev inequality is closely related to Polyak–Łojasiewicz inequality (or gradient dominance condition) and is commonly assumed in gradient-based systems to obtain convergence (for instance, see Kim et al. (2024)). This is formally stated in the following proposition and proved in Section E.3.
@@
where $\|\nabla_{\mathcal{M}}\mathcal{E}_{\lambda}(\theta,r)\|^{2}:=\|\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\|^{2}+\|\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]\|^{2}_{r_{t}}$. Moreover, if a log-Sobolev Inequality holds for a constant $\tau\in\mathbb{R}_{>0}$, i.e., for all $(\theta,r)\in\mathcal{M}$, we have
@@
Typically direct simulation of the gradient flow Eq. 7 is intractable as the derivative of the first variation of $\mathsf{R}^{\mathrm{E}}_{\lambda}$ involves $\nabla\log r_{t}$; instead, it is useful to identify the gradient flow with a McKean–Vlasov SDE, for which the gradient flow can be viewed as a Fokker–Planck equation, which can be simulated without access to this quantity. This SDE, which we term the PVI flow, is given by
$\mathrm{d}\theta_{t}=-\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\,\mathrm{d}t,\enskip\mathrm{d}Z_{t}=b(\theta_{t},r_{t},Z_{t})\,\mathrm{d}t+\sqrt{2\lambda_{r}}\,\mathrm{d}W_{t},\enskip\text{where }r_{t}=\mathrm{Law}(Z_{t}),$ (10)
where the drift is $b(\theta,r,\cdot):=-\nabla_{z}\delta_{r}\mathcal{E}[\theta,r]+\lambda_{r}\nabla_{z}\log p_{0}$ (with the first variation is given in Section E.3) and $W_{t}$ is a $d_{z}$-dimensional Wiener process. A connection between the Langevin diffusion, i.e., $\mathrm{d}Z_{t}=\nabla_{z}\log p(Z_{t},y)\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}$, and PVI flow can be observed with the fixed kernel $k{{}_{\theta}}(dx|z)=\delta_{z}(dx)$ and $\lambda_{r}=0$, namely, they both satisfy the same Fokker–Planck equation.
@@ 3.3 A practical algorithm
Input: initialization $(\theta_{0},\{Z_{0}\}_{i=1}^{M})$; regularization parameters $\{\lambda_{\theta},\lambda_{r}\}$; step-sizes $h_{\theta}$ and $h_{r}$; number of Monte Carlo samples $L$ (in Eq. 13); and preconditioner $\Psi=(\Psi^{\theta},\Psi^{r})$.
@@
     $r_{k-1}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\delta_{{Z}_{k-1,m}}$
     $\theta_{k}\leftarrow\theta_{k-1}-h_{\theta}\Psi^{\theta}\widehat{\nabla}_{\theta}\mathcal{E}_{\lambda}(\theta_{k-1},r_{k-1})$triangleright See Eq. 13
     $\hat{b}_{k}\leftarrow Z\mapsto-\widehat{\nabla}_{z}\delta_{r}\mathcal{E}[\theta_{k},r_{k-1}](Z)+\lambda_{r}\nabla_{z}\log p_{0}(Z)$triangleright See Eq. 13
@@
         $Z_{k,m}\leftarrow Z_{k-1,m}+h_{r}\Psi^{r}\hat{b}(Z_{k-1,m})+\sqrt{\lambda_{r}h_{r}\Psi^{r}}\eta_{k,m}$triangleright $\eta_{k,m}\sim\mathcal{N}(0,I_{d_{z}})$
@@
To produce a practical algorithm, we are faced with several practical issues. The first issue we tackle is the computation of gradients of expectations for which using standard automatic differentiation is insufficient. The second problem is that these gradients are often ill-conditioned and have different scales in each dimension. This is tackled using preconditioning resulting in adaptive stepsize. Finally, to produce computationally feasible algorithms, we show how to discretizate the PVI flow in both space and time. PVI can be found in Algorithm 1.
Computing the gradients. In the PVI flow, both the drift of the ODE and SDE include a gradient of an expectation with respect to parameters that define the distribution that is being integrated. Specifically, the terms that contain these gradients are $\nabla_{\theta}\mathcal{E}_{\lambda}$ and $\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}$. Fortunately, these gradients can be rewritten as an expectation as described in Prop. 6 for which the argument can be found in Section G.1.
@@
$\nabla_{z}\delta_{r}\mathcal{E}[\theta,r](z)=$ $\mathbb{E}_{p_{k}(\epsilon)}\left[(\nabla_{z}\phi_{\theta}\cdot[s_{\theta,r}-s_{p}])(z,\epsilon)\right].$ (12)
@@
^$\widehat{\nabla}_{\theta}\mathcal{E}(\theta,r):=\frac{1}{L}\sum_{l=1}^{L}\mathbb{E}_{r}[(\nabla_{z}\phi_{\theta}\cdot[s_{\theta,r}-s_{p}])(z,\epsilon_{l})],\enskip\widehat{\nabla}_{z}\delta_{r}\mathcal{E}[\theta,r]:=\frac{1}{L}\sum_{l=1}^{L}(\nabla_{z}\phi_{\theta}\cdot[s_{\theta,r}-s_{p}])(\cdot,\epsilon_{l}),$ (13)
where $\{\epsilon_{l}\}_{l=1}^{L}\overset{i.i.d.}{\sim}p_{k}$. This is an instance of a path-wise Monte-Carlo gradient estimator; a performant estimator that has been shown empirically to exhibit lower variance than other standard estimators (Kingma and Welling, 2014; Roeder et al., 2017; Mohamed et al., 2020).
Adaptive Stepsizes. One of the complexities of training neural networks is that their gradient is often poorly conditioned. As a result, for certain problems, the gradients computed from Eq. 11 and Eq. 12 can often produce unstable algorithms without careful tuning of the step sizes. We found it necessary to utilize preconditioners (Staib et al., 2019) to avoid this issue. Let $\Psi^{\theta}:\Theta\mapsto\mathbb{R}^{d_{\theta}\times d_{\theta}}$ and $\Psi^{r}:\mathbb{R}^{d_{z}}\mapsto\mathbb{R}^{d_{z}\times d_{z}}$ be the precondition for components $\theta$ and $r$ respectively, then the resulting preconditioned gradient flow is given by
$\mathrm{d}\theta_{t}=\Psi^{\theta}\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t}),\enskip\partial_{t}r_{t}=\nabla_{z}\cdot(r_{t}\Psi^{r}\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]).$ (14)
@@
$\mathrm{d}\theta_{t}=\Psi^{\theta}(\theta_{t})\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t}),\mathrm{d}Z_{t}=[\Psi^{r}(Z_{t})b(\theta_{t},r_{t},Z_{t})+\nabla_{z}\cdot\Psi^{r}(Z_{t})]\mathrm{d}t+\sqrt{2\lambda\Psi^{r}(Z_{t})}\mathrm{d}W_{t}.$ (15)
where $(\nabla\cdot\Psi^{r})_{i}=\sum_{j=1}^{d_{z}}\partial_{z_{j}}[(\Psi^{r})_{ij}]$ and $r_{t}=\mathrm{Law}(Z_{t})$. The equivalence between Eq. 14 and Eq. 15 is shown in Section G.2. A simple example for the preconditioner allows the $\theta$ and $\mathcal{P}(\mathbb{R}^{d_{z}})$ to have different time scales; ultimately, this results in different step sizes. Another more complex example of preconditioner $\Psi^{\theta}$ is the RMSProp (Tieleman and Hinton, 2012), and $\Psi^{r}$ we utilize a preconditioner inspired by RMSProp (see Section G.2). As with other related works (e.g., see (Li et al., 2016)), we found that the additional term $\nabla\cdot\Psi^{r}$ can be omitted in practice: it has little effect but incurs a large computational cost.
Discretization in both space and time. To obtain an actionable algorithm, we need to discretize the PVI flow in both space and time. For the space discretization, we propose to use a particle approximation for $r_{t}$, i.e., for a set of particles $\{Z_{t,m}\}_{m=1}^{M}$ with each satisfying $\mathrm{Law}(Z_{t,m})=r_{t}$, we utilize the approximation $r^{M}_{t}:=\frac{1}{M}\sum_{m=1}^{M}\delta_{Z_{t,m}}$ which converges almost surely to $r_{t}$ in the weak topology as $M\rightarrow\infty$ by the strong law of large numbers and a countable determining class argument (see, e.g. Theorem 1.1 of Schmon et al. (2020)). This approximation is key to making the intractable tractable, e.g., Eq. 1 is approximated by $q_{\theta,r^{M}_{t}}=\frac{1}{M}\sum_{m=1}^{M}k{{}_{\theta}}(x|Z_{t,m})$. One obtains a particle approximation to the PVI flow from the following ODE–SDE:
@@
where $[M]:=\{1,\ldots,M\}$. As for the time discretization, we employ Euler-Maruyama discretization with step-size $h$ which (using an appropriately defined preconditioner) can be decoupled into different stepsizes for $\theta$ and $Z$ components denoted by $h_{\theta}$ and $h_{r}$ respectively.

### v3

@@ 3.1 Free energy functional
$\operatorname*{arg\,min}_{(\theta,r)\in\mathcal{M}}\mathcal{E}(\theta,r),\quad\text{where}\enskip\mathcal{E}:\mathcal{M}\rightarrow\mathbb{R}:(\theta,r)\mapsto\int q_{\theta,r}(x)\log\frac{q_{\theta,r}(x)}{p(x,y)}\,\mathrm{d}x.$ (3)
@@
Assume that the evidence is bounded $\log p(y)<\infty$ and $k$ is bounded; then we have that $\mathcal{E}$ is (i) lower bounded, (ii) lower semi-continuous (l.s.c.), and (iii) non-coercive.
The proof can be found in Section E.1. Prop. 2 tells us that even though $\mathcal{E}$ possesses many of the properties one looks for in a meaningful minimization functional, it lacks coercivity (in the sense of Dal Maso, (2012, Definition 1.12)): a sufficient property to establish the existence of solutions. The key to showing non-coercivity is that we can construct a kernel $k{{}_{\theta}}(x|z)$ that does not depend on $z$. At first glance, this issue might seem contrived but we note that this problem is closely related to the problem of posterior collapse (Lucas et al.,, 2019; Wang et al.,, 2021). To address non-coercivity, we propose to utilize regularization and define the regularized free energy as:
@@
Under the assumptions of Prop. 2, if $\mathsf{R}_{\lambda}$ is coercive and l.s.c., then the regularized free energy $\mathcal{E}_{\lambda}$ is (i) lower bounded, (ii) l.s.c., (iii) coercive. Hence it admits at least one minimizer in $\mathcal{M}$.
The proof can be found in Section E.2. From here forward, we shall focus on regularizers of the form $\mathsf{R}^{\mathrm{E}}_{\lambda}:(\theta,r)\mapsto\lambda_{r}\mathsf{KL}(r,p_{0})+\lambda_{\theta}\mathsf{R}_{\theta}(\theta)$ where $\lambda=\{\lambda_{r},\lambda_{\theta}\}$ are the regularization parameters and $p_{0}$ is a predefined reference distribution. As long as $\mathsf{R}_{\theta}$ is l.s.c., coercive and $\lambda_{\theta},\lambda_{r}>0$, the resulting regularizer $\mathsf{R}^{\mathrm{E}}_{\lambda}$ will also be l.s.c. and coercive. There are many possible choices for $p_{0}$ and $\mathsf{R}_{\theta}$. For $p_{0}$, this regularizes solutions of the gradient flow toward it; as such, in settings where there is some knowledge or preference about $r$ at hand, we can set $p_{0}$ to reflect that. In our experiments, we utilize $p_{0}=\mathcal{N}(0,M)$ where $M$ is a positive definite (p.d.) matrix. As for $\mathsf{R}_{\theta}$, there are also many choices. In the context of neural networks, one natural choice is Tikhonov regularization $\frac{1}{2}\|\cdot\|^{2}$, resulting in weight decay (Hanson and Pratt,, 1988) for gradient descent (Loshchilov and Hutter,, 2019) which is a popular method for regularizing neural networks. In our experiments, we either use Tikhonov regularization or its simple variant $\|\theta\|^{2}_{M}:=\left\langle\theta,M\theta\right\rangle$.
@@ 3.2 Gradient flow
To solve the problem in Eq. 3, we construct a gradient flow that minimizes $\mathcal{E}_{\lambda}$. To this end, we endow the space $\mathcal{M}$ with a suitable notion of gradient $\nabla_{\mathcal{M}}\mathcal{E}_{\lambda}(\theta,r):=(\nabla_{\theta}\mathcal{E}_{\lambda},\nabla_{r}\mathcal{E}_{\lambda})$ where $\nabla_{\theta}$ and $\nabla_{r}$ denotes the Euclidean gradient and Wasserstein-$2$ gradient (Jordan et al.,, 1998), respectively. The latter gradient is given by $\nabla_{r}\mathcal{E}_{\lambda}(\theta,r):=-\nabla_{z}\cdot\left(r\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta,r\right]),$ where $\nabla_{z}\cdot$ denotes the standard divergence operator and $\delta_{r}$ denotes the first variation which is characterized in the following proposition.
@@
$(\dot{\theta}_{t},\dot{r}_{t})=-\nabla_{\mathcal{M}}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\iff\begin{array}[]{l}\dot{\theta}_{t}=-\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\\ \dot{r}_{t}=-\nabla_{r}\mathcal{E}_{\lambda}(\theta_{t},r_{t})=\nabla_{z}\cdot\left(r_{t}\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]\right).\end{array}$ (7)
We now establish that the above gradient flow dynamic is contractive and that if a log-Sobolev inequality Eq. 9 holds, one can also establish exponential convergence. The log-Sobolev inequality is closely related to Polyak–Łojasiewicz inequality (or gradient dominance condition) and is commonly assumed in gradient-based systems to obtain convergence (for instance, see Kim et al., (2024)). This is formally stated in the following proposition and proved in Section E.3.
@@
where $\|\nabla_{\mathcal{M}}\mathcal{E}_{\lambda}(\theta,r)\|^{2}:=\|\nabla_{\theta}\mathcal{E}_{\lambda}(\theta,r)\|^{2}+\|\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta,r]\|^{2}_{r}$. Moreover, if a log-Sobolev Inequality holds for a constant $\tau\in\mathbb{R}_{>0}$, i.e., for all $(\theta,r)\in\mathcal{M}$, we have
@@
Typically direct simulation of the gradient flow Eq. 7 is intractable as the derivative of the first variation of $\mathsf{R}^{\mathrm{E}}_{\lambda}$ involves $\nabla_{z}\log r_{t}$; instead, it is useful to identify the gradient flow with a McKean–Vlasov SDE, for which they share the same Fokker–Planck equation. The key distinction is that the SDE can be simulated without access to $\nabla_{z}\log r_{t}$. This SDE, which we term the PVI flow, is given by
$\mathrm{d}\theta_{t}=-\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\,\mathrm{d}t,\enskip\mathrm{d}Z_{t}=b(\theta_{t},r_{t},Z_{t})\,\mathrm{d}t+\sqrt{2\lambda_{r}}\,\mathrm{d}W_{t},$ (10)
where $r_{t}=\mathrm{Law}(Z_{t})$, the drift is $b(\theta,r,\cdot):=-\nabla_{z}\delta_{r}\mathcal{E}[\theta,r]+\lambda_{r}\nabla_{z}\log p_{0}$ (with the first variation given in Prop. 4) and $W_{t}$ is a $d_{z}$-dimensional Wiener process. A connection between the Langevin diffusion, i.e., $\mathrm{d}Z_{t}=\nabla_{z}\log p(Z_{t},y)\,\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}$, and PVI flow can be observed with the fixed kernel $k{{}_{\theta}}(\mathrm{d}x|z)=\delta_{z}(\mathrm{d}x)$ and $\lambda_{r}=0$, namely, they both satisfy the same Fokker–Planck equation.
@@ 3.3 A practical algorithm
Input: initialization $(\theta_{0},\{Z_{0,m}\}_{m=1}^{M})$; regularization parameters $\{\lambda_{\theta},\lambda_{r}\}$; step-sizes $h_{\theta}$ and $h_{r}$; number of Monte Carlo samples $L$ (for Eqs. 13 and 14); and preconditioner $\Psi=(\Psi^{\theta},\Psi^{r})$.
@@
     $r^{M}_{k-1}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\delta_{{Z}_{k-1,m}}$
     $\theta_{k}\leftarrow\theta_{k-1}-h_{\theta}\Psi^{\theta}\widehat{\nabla}_{\theta}\mathcal{E}_{\lambda}(\theta_{k-1},r^{M}_{k-1})$triangleright See Eq. 13
     $\hat{b}_{k}\leftarrow Z\mapsto-\widehat{\nabla}_{z}\delta_{r}\mathcal{E}[\theta_{k},r^{M}_{k-1}](Z)+\lambda_{r}\nabla_{z}\log p_{0}(Z)$triangleright See Eq. 14
@@
         $Z_{k,m}\leftarrow Z_{k-1,m}+h_{r}\Psi^{r}\hat{b}_{k}(Z_{k-1,m})+\sqrt{\lambda_{r}h_{r}\Psi^{r}}\eta_{k,m}$triangleright $\eta_{k,m}\sim\mathcal{N}(0,I_{d_{z}})$
@@
To produce a practical algorithm, we are faced with several practical issues. The first issue we tackle is the computation of gradients of expectations for which using standard automatic differentiation is insufficient. The second problem is that these gradients are often ill-conditioned and have different scales in each dimension. This is tackled using preconditioning resulting in adaptive stepsize. Finally, to produce computationally feasible algorithms, we show how to discretize the PVI flow in both space and time. PVI is summarised in Algorithm 1.
Computing the gradients. In the PVI flow, both the drift of the ODE and SDE include a gradient of an expectation with respect to parameters that define the distribution that is being integrated. Specifically, the terms that contain these gradients are $\nabla_{\theta}\mathcal{E}_{\lambda}$ and $\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}$. Fortunately, these gradients can be rewritten as an expectation (as described in Prop. 6) for which the parameters being differentiated w.r.t. is only found in the integrand (see derivation in Section G.1).
@@
$\nabla_{z}\delta_{r}\mathcal{E}[\theta,r](z)=$ $\mathbb{E}_{p_{k}(\epsilon)}\left[(\nabla_{z}\phi_{\theta}\cdot[s_{\theta,r}-s_{p}])(z,\epsilon)\right],$ (12)
@@
^$\widehat{\nabla}_{\theta}\mathcal{E}(\theta,r)$ $:=\frac{1}{L}\sum_{l=1}^{L}\mathbb{E}_{z\sim r}[(\nabla_{z}\phi_{\theta}\cdot[s_{\theta,r}-s_{p}])(z,\epsilon_{l})],$ (13)
^$\widehat{\nabla}_{z}\delta_{r}\mathcal{E}[\theta,r]$ $:=\frac{1}{L}\sum_{l=1}^{L}(\nabla_{z}\phi_{\theta}\cdot[s_{\theta,r}-s_{p}])(\cdot,\epsilon_{l}),$ (14)
where $\{\epsilon_{l}\}_{l=1}^{L}\overset{i.i.d.}{\sim}p_{k}$. This is an instance of a path-wise Monte-Carlo gradient estimator; a performant estimator that has been shown empirically to exhibit lower variance than other standard estimators (Kingma and Welling,, 2014; Roeder et al.,, 2017; Mohamed et al.,, 2020).
Adaptive Stepsizes. One of the complexities of training neural networks is that their gradient is often poorly conditioned. As a result, for certain problems, the gradients computed from Eq. 11 and Eq. 12 can often produce unstable algorithms without careful tuning of the step sizes. When this occurs, we utilize preconditioners (Staib et al.,, 2019) to avoid this issue. Let $\Psi^{\theta}:\Theta\mapsto\mathbb{R}^{d_{\theta}\times d_{\theta}}$ and $\Psi^{r}:\mathbb{R}^{d_{z}}\mapsto\mathbb{R}^{d_{z}\times d_{z}}$ be the precondition for components $\theta$ and $r$ respectively, then the resulting preconditioned gradient flow is given by
$\mathrm{d}\theta_{t}=-\Psi^{\theta}\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\,\mathrm{d}t,\enskip\partial_{t}r_{t}=\nabla_{z}\cdot(r_{t}\Psi^{r}\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]).$ (15)
@@
$\mathrm{d}\theta_{t}$ $=-\Psi^{\theta}(\theta_{t})\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\,\mathrm{d}t,$ (16)
$\mathrm{d}Z_{t}$ $=[\Psi^{r}(Z_{t})b(\theta_{t},r_{t},Z_{t})+\nabla_{z}\cdot\Psi^{r}(Z_{t})]\,\mathrm{d}t+\sqrt{2\lambda\Psi^{r}(Z_{t})}\mathrm{d}W_{t},$ (17)
where $(\nabla_{z}\cdot\Psi^{r})_{i}=\sum_{j=1}^{d_{z}}\partial_{z_{j}}[(\Psi^{r})_{ij}]$ and $r_{t}=\mathrm{Law}(Z_{t})$. The equivalence between Eq. 15 and Eqs. 16 and 17 is shown in Section G.2. A simple example for the preconditioner allows the $\theta_{t}$ and $Z_{t}$ to have different time scales; ultimately, this results in different step sizes. Another more complex example of preconditioner $\Psi^{\theta}$ is the RMSProp (Tieleman and Hinton,, 2012), and $\Psi^{r}$ we utilize a preconditioner inspired by RMSProp (see Section G.2). As with other related works (e.g., see Li et al., (2016)), we found that the additional term $\nabla_{z}\cdot\Psi^{r}$ can be omitted in practice: it has little effect but incurs a large computational cost.
Discretization in both space and time. To obtain an actionable algorithm, we need to discretize the PVI flow in both space and time. For the space discretization, we propose to use a particle approximation for $r_{t}$, i.e., for a set of particles $\{Z_{t,m}\}_{m=1}^{M}$ with each satisfying $\mathrm{Law}(Z_{t,m})=r_{t}$, we utilize the approximation $r^{M}_{t}:=\frac{1}{M}\sum_{m=1}^{M}\delta_{Z_{t,m}}$ which converges almost surely to $r_{t}$ in the weak topology as $M\rightarrow\infty$ by the strong law of large numbers and a countable determining class argument (e.g., see Schmon et al., (2020, Theorem 1.1)). This approximation is key to making the intractable tractable, e.g., Eq. 1 is approximated by $q_{\theta,r^{M}_{t}}=\frac{1}{M}\sum_{m=1}^{M}k{{}_{\theta}}(x|Z_{t,m})$. One obtains a particle approximation to the PVI flow from the following ODE–SDE:
@@
where $[M]:=\{1,\ldots,M\}$. As for the time discretization, we employ Euler–Maruyama discretization with step-size $h$ which (using an appropriately defined preconditioner) can be decoupled into different stepsizes for $\theta_{t}$ and $Z_{t}$ denoted by $h_{\theta}$ and $h_{r}$ respectively.